OpenStudy (anonymous):

Check my answer: The kitchen in Shane’s house is a square. The points P(0, 7) and Q(4, 1) represent the opposite vertices of his kitchen in a drawing on a coordinate grid. Which of these ordered pairs could be a third vertex? (5, 7) (-1, 1) (-1, 2) (4, 7) I got (-1,2)

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

@amistre64

mathslover (mathslover):

Its (-1,2) http://openstudy.com/study#/updates/4f47c174e4b00aeb3e3ef2db

OpenStudy (amistre64):

? 41 24 (midpoint) 07 ? this looks rather squarish to me; not as , hmmm

mathslover (mathslover):

Let the required point be : \(x,y\) then : \(\sqrt{(x-0)^2 + (y-7)^2 }= \sqrt{(x-4)^2 + (y-1)^2 }\) \(x^2 + y^2 + 49 - 14y = x^2 + 16 - 8x + y^2 + 1 - 2y \) \(49 - 14y + 8x - 16 -1 + 2y = 0\) \(32 - 12y + 8x = 0 \) \(8x - 12y = 32\) \(2x - 3y = 8 \) From the options, only (-1,2) satisfies the given relationship.

mathslover (mathslover):

Oopps, I meant 8x - 12y = -32 and 2x -3y = -8

OpenStudy (amistre64):

-1,2 is not going to be it ... in my view

OpenStudy (amistre64):

but then i am blind most of the time

OpenStudy (amistre64):

center a circle at 2,4 (x-2)^2 + (y-4)^2 = r^2 0 7 --------------------- 4 + 9 = 13 (x-2)^2 + (y-4)^2 = 15 -1 2 --------------------- 9 + 4 yep

OpenStudy (anonymous):

so whats the answer...? @amistre64

OpenStudy (amistre64):

i was checking to see if -1 2 worked in my circle idea, and it did ... , so id go with -1,2

OpenStudy (anonymous):

@amistre64 oh alrighty thanks guys :)

OpenStudy (amistre64):

if we try the other options .... since not all points on a circle are squares .... (x-2)^2 + (y-4)^2 = 13 5 7 -------------------- 9 + 9 not it (x-2)^2 + (y-4)^2 = 13 -1 1 -------------------- 9 + 9 not it (-1, 2) , we know this is fine (x-2)^2 + (y-4)^2 = 13 4 7 ------------------- 4 + 9 = 13 this one would give me pause to consider

OpenStudy (amistre64):

(0, 7) (4, 1) -7 -7 ----------- 4,-6 ; slope is -6/4 = -3/2 ; perp is 2/3 y-4 = 2/3(x-2) y-4 = 2/3(x-2) 7 4 -------------- 3 = 4/3 not it y-4 = 2/3(x-2) 2 -1 -------------- -2 = -2 , its -1,2 fer sure